OpenStudy (anonymous):

What values best approximate the solutions of x^2 + 4x - 3.5 = 0?

OpenStudy (anonymous):

help?

OpenStudy (pfenn1):

You can use the quadratic formula to solve this equation.

OpenStudy (anonymous):

ok I think its -4.8 and 0.8 .. am I correct?

OpenStudy (pfenn1):

I got a different answer but let me recalculate I got \[\left( \left( -4\pm \sqrt{2} \right)\over 2 \right)=-2\pm \left( \sqrt{2}\over 2 \right)\]

OpenStudy (pfenn1):

or about x=-2.7, x=-1.2 I hope I didn't make any mistakes

OpenStudy (anonymous):

so no real solutions Idk I don't see that :/ -4.8 and -3.5 - 4.8 and 0.8 -3.5 and 0.8 no real solutions

OpenStudy (anonymous):

help? :/

OpenStudy (pfenn1):

Sorry, I had to go make dinner for the family. I don't know where my mistake was but if I plug -4.8 and 0.8 into the equation, they are approximately correct, so I would go with that.